i need help with all parts. i will rate.

thank you very much.f(x,y) = xy on the line 4x+9y=36 using Lagrange Multipliers, the value of is 2 O 2 3 O2If the position vector is r(t) = (cost, sin t, t) the velocity and acceleration are orthogonal only at time O t=0 O t=1 O t=TThe vector < 10, 12,4> is normal to the surface x2+y3+24 =18 at which point? O (0,0,0) O (4,1,1) O (2,3,-4) O (5,-2, 1)
Answer #1

More the objective is to maximize or minimise the function fm,y). So fu,y) is called djective function constraints restrictsMere, objective feinetion f(wy)=xy and constrainst id untay=36 from Lagrangs multiplier, of de let gcn,y) = untay-3620 » lets>79-6914+) i + (o+copy); 79 ܪ ܢܐ + q j Twhile partially derivating with a y remains as constant as vice versas e kan= knany ;2.11a 2 2) Criven position vector r(t) = {cost, sint, t²). we know, velocity w(t) = f(t) and acceleration act) = vct) so coorthogonal of ture vectors a and ħ are Then a. To =0 so, consider VCE) and act). TCE). ä(t) = 0 > (sintitest j+2+$.fcesti-sin18. 3) Given vector (10, 12,4%. is normal to surface nity² + 24 =18 we know, Gradient of a vector is Normal to the swface ThePartially derivate & with a to & = stad az fz f (x² + y² +2418) cot uz? fz =u23 : orthogonal vector is (fon, fy, fz) 3y?, 423